Suppose that Total revenue of the firm is
TR = 5Q + 3Q2 and TC is TC = Q + 4Q2.
Obtain the quantity that maximizes profit and hence maximum profit.
Suppose that Total revenue of the firm is
TR = 5Q + 3Q2 and TC is TC = Q + 4Q2.
Obtain the quantity that maximizes profit and hence maximum profit.
1
Expert's answer
2019-11-06T11:00:57-0500

Solution:


"5+6Q=1+8Q"


"MC=TC'=1+4\times2Q=1+8Q"

"MR=TR'=5+3\times2Q=5+6Q"


"2Q=4"


"Q=2""Pr=TR-TC"


"Pr=5Q+3Q^2-Q-4Q^2=4Q-Q^2"

"Pr=4\times2-4"


"Pr=4"

Answer Q=2, profit = 4